OpenStudy (itrymath):

Solve for x: -3|2x + 6| = -12

OpenStudy (itrymath):

i think they are bot -1

OpenStudy (itrymath):

x = -9 and x = 3 No solution x = 1 and x = 5 x = -1 and x = -5 those are answer choices

Awolflover1 (awolflover1):

\[−3(|2x+6|)=−12\]

OpenStudy (itrymath):

how

Awolflover1 (awolflover1):

Divide both sides by -3

OpenStudy (itrymath):

?

OpenStudy (itrymath):

why

OpenStudy (itrymath):

that equals 4 i think

Awolflover1 (awolflover1):

to get your answer :) let's go step by step ^_^

Awolflover1 (awolflover1):

yes correct

OpenStudy (itrymath):

and....

OpenStudy (itrymath):

whats that 4 do lol

Awolflover1 (awolflover1):

now \[|2x+6|=4\]

Awolflover1 (awolflover1):

\[2x+6−6=4−6\] Subtract 6 from both sides

OpenStudy (itrymath):

OMG

OpenStudy (itrymath):

YOU DA BEST

OpenStudy (itrymath):

I understand like 10000000% now

OpenStudy (itrymath):

Never giving up on math even when it gets hard ;)

Awolflover1 (awolflover1):

:)

Awolflover1 (awolflover1):

\[4-6=? \]

OpenStudy (itrymath):

-2

OpenStudy (itrymath):

why'd you ask ?

Awolflover1 (awolflover1):

yes \[2x=−2\] Divide both sides by 2

OpenStudy (itrymath):

lmao i meant i did the whole think already

OpenStudy (itrymath):

i know the answer xP

Awolflover1 (awolflover1):

Awesome! How about the other?

OpenStudy (itrymath):

x=-1 x=5

Awolflover1 (awolflover1):

Very Close but x=5 is incorrect

OpenStudy (itrymath):

WHAT

OpenStudy (itrymath):

HOW

OpenStudy (itrymath):

I did the work!

OpenStudy (itrymath):

you're lieing

Awolflover1 (awolflover1):

\[2x+6−6=−4−6\] \[2x=−10\] it is correct but not positive ;)

OpenStudy (itrymath):

impossible

OpenStudy (itrymath):

your're wrong

OpenStudy (itrymath):

omg

Awolflover1 (awolflover1):

Noo! try it again! \[−4−6\]

OpenStudy (itrymath):

its 2x-6=4 then +6 plus 4 euals 10 then 10 divide 2 equals 5!

OpenStudy (itrymath):

>:(

OpenStudy (itrymath):

:c

Awolflover1 (awolflover1):

2x=−10 Divide both sides by 2

Awolflover1 (awolflover1):

you're very close!

OpenStudy (itrymath):

look

OpenStudy (itrymath):

those lines on the side of quation mean 2x+6 and 2x-6

OpenStudy (itrymath):

!!!!!!!!!

Awolflover1 (awolflover1):

let's do this again You're Correct but its not Positive!

Awolflover1 (awolflover1):

\[-10\div2=?\]

OpenStudy (itrymath):

-5

Awolflover1 (awolflover1):

there you go!

OpenStudy (itrymath):

nooooo but how

OpenStudy (itrymath):

why isnt it +6 to the 4

Awolflover1 (awolflover1):

\[2x+6=−4\]

OpenStudy (itrymath):

noonononono

OpenStudy (itrymath):

impossible

OpenStudy (itrymath):

its the bar things arent around the 4

Awolflover1 (awolflover1):

@563blackghost correct me please im confused

OpenStudy (itrymath):

OOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (itrymath):

im so sorry

OpenStudy (itrymath):

omg

OpenStudy (itrymath):

lol

OpenStudy (itrymath):

i figured it out

Awolflover1 (awolflover1):

No! Its okay! Im confused!

OpenStudy (itrymath):

no you're correct

OpenStudy (itrymath):

sorry

Awolflover1 (awolflover1):

haha its okay we all have are mistakes

OpenStudy (itrymath):

omg i looked back i forgot my original equation was -3|2x + 6| = -12

OpenStudy (itrymath):

can you help me with another one but ill do work this time and you just tell me if im right or wrong

Awolflover1 (awolflover1):

alright (:

OpenStudy (itrymath):

Solve for x: |x| - 8 = -5 (2 points) x = 3 and x = -3 x = 3 and x = 13 x = -13 and x = -3 No solution

OpenStudy (itrymath):

yeah the reason i dont know dis is because i didnt go to school

OpenStudy (itrymath):

this*

Awolflover1 (awolflover1):

haha that's okay ^_^ Well it can't be No Solution because there is a solution so we can cross that out

OpenStudy (itrymath):

no i mean thats why im on this website lmao im doing this by myself no tips

Awolflover1 (awolflover1):

Well that's what we are here for to teach people (:

Awolflover1 (awolflover1):

I'll give you a hint |x|−8+8=−5+8 \[-5+8=? \]

OpenStudy (itrymath):

x=3 and x=13

OpenStudy (itrymath):

and if you dont mind i have 2 more and ill be honest this is my homework but im doing it myself because i want to know how to do this and also i really appreciate your time i know im annoying and i push it wish there was a way i can repay you lol

Awolflover1 (awolflover1):

x=3 is correct but x=13 is incorrect and Its alright haha

OpenStudy (itrymath):

O My GoD

OpenStudy (itrymath):

my heads going to explode lol :)

OpenStudy (itrymath):

x=-3

Awolflover1 (awolflover1):

yes there we go!

OpenStudy (itrymath):

Solve for x: -4|x + 5| = -16 (2 points) x = -1, x = 9 x = -1, x = -9 No solution x = , x = -

OpenStudy (itrymath):

no tips

OpenStudy (itrymath):

actually yes tips i have no idea hehe

Awolflover1 (awolflover1):

x+5−5=4−5 \[4-5=? \]

OpenStudy (itrymath):

waoit -4 divide -16 correct? will be 4?

OpenStudy (itrymath):

yay im correct

Awolflover1 (awolflover1):

haha yes

OpenStudy (itrymath):

x=-1?

OpenStudy (itrymath):

or 0

Awolflover1 (awolflover1):

yep =) x=-1

Awolflover1 (awolflover1):

\[x+5−5=−4−5\] now -4-5=?

OpenStudy (itrymath):

x=1

OpenStudy (itrymath):

oh sorry -1

Awolflover1 (awolflover1):

x=-1 correct the other one ? x= ?

OpenStudy (itrymath):

wait but why does the sign change instead of the total

OpenStudy (itrymath):

-1